OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (texaschic101):

3 | x - 8 | + 15 = 18 --- subtract 15 from both sides 3 | x - 8 | = 3 --- divide both sides by 3 | x - 8 | = 1 (x - 8) = 1 x = 1 + 8 x = 9 -(x - 8) = 1 -x + 8 = 1 -x = -7 x = 7 x = 9 and x = 7

OpenStudy (anonymous):

@texaschic101 knows his stuff. :)